OpenStudy (anonymous):

Aneesha travels at a rate of 50 miles per hour. Morris is traveling 3 feet per second less than Aneesha. Which is the best estimate of the speed Morris is traveling? 1 mile = 5,280 feet 45 miles per hour 46 miles per hour 47 miles per hour 48 miles per hour

OpenStudy (welshfella):

use the following:- there are 5280 feet in a mile and 3600 seconds in an hour

OpenStudy (anonymous):

It's D right?

OpenStudy (welshfella):

50 mph = 50*5280 ft / hour = 50*5280 / 3600 ft / sec 73.33 ft/sec so Morrius speed is 70.433 ft/sec in mph this is (70.33* 3600 )/ 5280 = ?

OpenStudy (welshfella):

what do you get for that?

OpenStudy (welshfella):

47.95 mph

OpenStudy (welshfella):

D is correct

OpenStudy (anonymous):

Thank you!

OpenStudy (welshfella):

yw